A wire made of copper is found to have a change in resistance of 17% after its temperature change. If the resistance of this wire was 8.66 Ω at 20 0C, and the resistivity after the change in temperature is found to have decreased, determine the following: 

If the wire is connected to 14 V battery, how many electrons pass through a cross-sectional area A in 3 min after temperature change, assuming the wire is an ohmic conductor          


Select one:

 2.2×1021 electrons

1.6×1019 electrons 

3.6×1019 electrons

1.2×1021 electrons


1
Expert's answer
2021-11-25T11:51:41-0500
"q=Ne=It\\Ne=\frac{U}{R'}t\\(1.6\cdot10^{-19})N=\frac{14}{(1-0.17)8.66}(3\cdot60)\\N=2.2\cdot10^{21}"